OpenStudy (anonymous):

What happens when you try to find the slope and y intercept for the equation. 3(y+2) + 6x = 3(4 + 2x + y)

OpenStudy (anonymous):

\[ \ \ \ 3(y+2)+6x=3[4+2x+y]\\ \ \ \ 3y+6+6x=12+6x+3y\\ \ \ \ 6\ne12\\ \text{The equation has no solutions.} \]

OpenStudy (anonymous):

same thing i got but wasnt so sure thnx bro :)